Kabbed:

We look for the minimum of each function. For f (x) = 3x2 + 12x + 16: We derive the function: f '(x) = 6x + 12

Kabbed:

@Narad @Vocaloid

Kabbed:

@Mercury

Kabbed:

what is derive the function mean

Kabbed:

what does it mean when he derives the function help please just that simple question question what is derive

Narad:

When f'(x)=0 This means that we have a critical point. 6x+12=0 6x=-12 x=-12/6=-2 There is a minimum at (-2, 4)